# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v2

@@ 10.4 Proof of Proposition 2
where the first identity follows from the definition of composition and the second identity employs Toland’s duality (Toland,, 1978) to rewrite difference of convex functions optimization.
By using infimal convolutions (Rockafellar,, 1997, Theorem 16.4), we can reformulate $g^{*}$:
@@
where the first identity follows from the definition of the convex conjugate, the second identity introduces an additional variable to make this an equality-constrained optimization problem, the third identity takes the Lagrange dual (which is a strong dual since the problem maximizes a concave objective with a single equality constraint), the fourth identity rearranges the expressions, the fifth identity exploits unboundedness of $\bm{x}$, the sixth identity uses the definition of convex conjugates and the final identity replaces the feasible set $\theta\in\mathbb{R}$ with the domain of $L^{\star}$ without loss of generality as this is an $\inf$ problem.
@@
DC maximization terms similar to the one dealt by Lemma 1 appear on the left-hand side of the constraints of Inter-ARO (cf. formulation in Proposition 1). These constraints would admit a tractable reformulation for the case without auxiliary data because the $\inf$ term in the reformulation presented in Lemma 1 does not appear in such cases. To see this, eliminate the second norm (the one associated with auxiliary data) by taking $\widehat{\lambda}=0$, which will cause the constraint $\lvert\theta\rvert\cdot\lVert\bm{z}\rVert_{q^{\star}}\leq\widehat{\lambda}$ to force $\bm{z}=\mathbf{0}$, and the alternative formulation will thus be:
@@
where we used the fact that $L=L^{**}$ and $\sup_{\theta\in\mathrm{dom}(L)}\lvert\theta\rvert=\mathrm{Lip}(L)$ since $L$ is closed convex (Rockafellar,, 1997, Corollary 13.3.3). Hence, the DC maximization can be represented with a convex function with an additional convex inequality, making the constraints tractable for the case without auxiliary data. For the case with auxiliary data, however, the $\sup_{\theta}\inf_{\bm{z}}$ structure makes these constraints equivalent to two-stage robust constraints (with uncertain parameter $\theta$ and adjustable variable $\bm{z}$), bringing an adjustable robust optimization (Ben-Tal et al.,, 2004; Yanıkoğlu et al.,, 2019) perspective to Inter-ARO. By using the univariate representation $\ell_{\bm{\beta}}^{\alpha}(\bm{x},y)=L^{\alpha}(y\cdot\bm{\beta}^{\top}\bm{x})$Inter-ARO can be written as
@@
This reduces problem (41) to a generic two-stage robust optimization problem with random recourse (Subramanyam et al.,, 2020, Problem 1) which is proven to be NP-hard even if $S^{\alpha*}$ was constant (Guslitser,, 2002). ∎
10.5 Proof of Theorem 1
Consider the reformulation 38 of Inter-ARO that we introduced in the proof of Proposition 2. For any $i\in[N]$ and $j\in[\widehat{N}]$, the corresponding constraint in the first group of ‘adjustable robust’ ($\forall$,\ $\exists$ , ∃) constraints will be:
@@ 10.5 Proof of Theorem 1
Notice that this is a safe approximation, since any fixed $\bm{z}$ satisfying the latter system is a feasible static solution in the former system, meaning that for every realization of $\theta$ in the first system, the inner $\exists\bm{z}$ can always ‘play’ the same $\bm{z}$ that is feasible in the latter system (hence the latter is named the static relaxation, Bertsimas et al., 2015). In the relaxed system, we can drop $\forall\theta$ and keep its worst-case realization instead:
@@
The term $\sup_{\theta\in\mathrm{dom}(L^{*})}\{-L^{\alpha*}(\theta)+\theta\cdot\bm{\beta}^{\top}\bm{x}^{i}+\theta\cdot\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i})\}$ is the definition of the biconjugate $L^{\alpha**}(\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))$. Since $L^{\alpha}$ is a closed convex function, we have $L^{\alpha**}=L^{\alpha}$ (Rockafellar,, 1997, Corollary 12.2.1). Moreover, $\sup_{\theta\in\mathrm{dom}(L^{*})}\{\lvert\theta\rvert\}$ is an alternative representation of the Lipschitz constant of the function $L^{\alpha}$ (Rockafellar,, 1997, Corollary 13.3.3), which is equal to $1$ as we showed earlier. The adjustable robust constraint thus reduces to:
@@
as a result of the static relaxation. This relaxed reformulation applies to all $i\in[N]$ and $j\in[\widehat{N}]$ as well as to the second group of adjustable robust constraints analogously. Replacing each constraint of 38 with this system concludes the proof.∎
10.6 Proof of Corollary 2
@@
where we also used the definition of $L^{\alpha}$. This problem is identical to DR-ARO, which means that feasible solutions of DR-ARO are feasible for 16 if the additional variables $(\widehat{\lambda},\widehat{\bm{s}},\bm{z}^{l}_{ij})$ are set to zero, concluding the first statement of the corollary.
The second statement is immediate since $\widehat{\varepsilon}\rightarrow\infty$ forces $\widehat{\lambda}=0$ due to the term $\widehat{\varepsilon}\widehat{\lambda}$ in the objective of 16, and this proof shows in such a case 16 reduces to DR-ARO (which is identical to Inter-ARO when $\varepsilon\rightarrow\infty$ by definition). ∎
10.7 Proof of Observation 2
@@ 10.8 Proof of Proposition 3
and since $d$ is a feature-label metric (cf. Definition 1) the minimum is well-defined (Villani et al.,, 2009, Theorem 4.1). We name the optimal solutions to the above problem the optimal couplings. Let $\Pi^{1}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{1})$ and let $\Pi^{2}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{2})$ and define $\Pi^{c}=\lambda\cdot\Pi^{1}+(1-\lambda)\cdot\Pi^{2}$. We have
@@
10.9 Proof of Theorem 2
Since each result in the statement of this theorem is abridged, we will present these results sequentially as separate results. We review the existing literature to characterize $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, in a similar fashion with the results presented in (Selvi et al.,, 2022, Appendix A) for the logistic loss, by revising them to the adversarial loss whenever necessary. The $N$-fold product distribution of $\mathbb{P}^{0}$ from which the training set $\mathbb{P}_{N}$ is constructed is denoted below by $[\mathbb{P}^{0}]^{N}$.
@@ 10.9 Proof of Theorem 2
The statement follows from Theorem 18 of Kuhn et al., (2019). The presented decay rate $\mathcal{O}(N^{-1/n})$ of $\varepsilon$ as $N$ increases is optimal (Fournier and Guillin,, 2015). ∎
Now that we gave a confidence for the radius $\varepsilon$ of $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, we analyze the underlying optimization problems. Most of the theory is well-established for logistic loss function, and in the following we show that similar results follow for the adversarial loss function. For convenience, we state DR-ARO again by using the adversarial loss function as defined in Observation 1:
@@
holds for all $\eta\in(0,1)$ and all optimizers $\bm{\beta}^{\star}$ of 44.
@@
The statement follows from Theorem 19 of Kuhn et al., (2019) given that $\ell^{\alpha}_{\bm{\beta}}$ is a finite-valued continuous loss function. ∎
Theorem 5 states that the optimal value of 44 overestimates the true loss with arbitrarily high confidence $1-\eta$. Despite the desired overestimation of the true loss, we show that 44 is still asymptotically consistent if we restrict the set of admissible $\bm{\beta}$ to a bounded set [^2] Note that, this is without loss of generality given that we can normalize the decision boundary of linear classifiers..
@@
If we show that there exists $\bm{\xi}^{0}\in\Xi$ and $C>0$ such that $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)\leq C(1+d(\bm{\xi},\bm{\xi}^{0}))$ holds for all $\bm{\beta}\in\mathcal{H}$ and $\bm{\xi}\in\Xi$ (that is, the adversarial loss satisfies a growth condition), the statement will follow immediately from Theorem 20 of (Kuhn et al.,, 2019).
@@
$\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))\leq C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{0}]})$
@@
$C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{0}]})$ $=C(1+\lVert\bm{x}\rVert_{q}+\kappa\cdot\mathds{1}{[y\neq y^{0}]})$
@@
So far, we reviewed tight characterizations for $\varepsilon$ so that the ball $\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ includes the true distribution $\mathbb{P}^{0}$ with arbitrarily high confidence, proved that the DRO problem 44 overestimates the true loss, while converging to the true problem asymptotically as the confidence $1-\eta$ increases and the radius $\varepsilon$ decreases simultaneously. Finally, we discuss that for optimal solutions $\bm{\beta}^{\star}$ to 44, there are worst case distributions $\mathbb{Q}^{\star}\in\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ of nature’s problem that are supported on at most $N+1$ atoms.
@@
If we restrict the hypotheses $\bm{\beta}$ to a bounded set $\mathcal{H}\subseteq\mathbb{R}^{n}$, then there are distributions $\mathbb{Q}^{\star}\in\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ that are supported on at most $N+1$ atoms and satisfy:
@@
The proof follows from (Yue et al.,, 2022). ∎
See the proof of Selvi et al., (2022, Theorem 8) and the discussion that follows for insights and further analysis on these results presented.
10.10 Proof of Theorem 3
@@
$\widehat{\varepsilon}_{1}\geq\begin{cases}\left(\dfrac{\log(\widehat{c}_{1}/\eta_{2})}{\widehat{c}_{2}\cdot\widehat{N}}\right)^{1/\max\{n,2\}}&\text{if }\widehat{N}\geq\dfrac{\log(\widehat{c}_{1}/\eta_{2})}{\widehat{c}_{2}}\\ \left(\dfrac{\log(\widehat{c}_{1}/\eta_{2})}{\widehat{c}_{2}\cdot\widehat{N}}\right)^{1/\widehat{a}}&\text{otherwise}\end{cases}$
@@
11 EXPONENTIAL CONIC REFORMULATION OF DR-ARO
@@
Following the conic modeling guidelines of MOSEK ApS, (2023), for new variables $v^{+}_{i},w^{+}_{i}\in\mathbb{R}$, the first constraint can be written as
@@
12 FURTHER DETAILS ON NUMERICAL EXPERIMENTS
12.1 UCI Experiments
Preprocessing UCI datasets
We experiment on 10 UCI datasets (Kelly et al.,, 2023) (cf. Table 3). We use Python 3 for preprocessing these datasets. Classification problems with more than two classes are converted to binary classification problems (most frequent class/others). For all datasets, numerical features are standardized, the ordinal categorical features are left as they are, and the nominal categorical features are processed via one-hot encoding. As mentioned in the main paper, we obtain auxiliary (synthetic) datasets via SDV, which is also implemented in Python 3.
@@ 12.1 UCI Experiments
Detailed misclassification results on the UCI datasets
Table 4 contains detailed results on the out-of-sample error rates of each method on 10 UCI datasets for classification. All parameters are $5$-fold cross-validated: Wasserstein radii from the grid $\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0,1,2,5,10\}$ ($1010^{-6},10^{-5},2,5,$ 10 are rarely selected, but we did not change our grid in order not to introduce a bias), $\kappa$ from the grid $\{1,\sqrt{n},n\}$ the weight parameter of ARO+Aux from grid $\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0,1\}$. We fix the norm defining the feature-label metric to the $\ell_{1}$-norm, and test $\ell_{2}$-attacks, but other choices with analogous results are also implemented.
@@
Finally, we demonstrate that our theory, especially DRO+ARO+Aux, contributes to the DRO literature even without adversarial attacks. In this case of $\alpha=0$, ERM and ARO would be equivalent, and DRO+ARO would reduce to the traditional DR LR model (Shafieezadeh-Abadeh et al.,, 2015). ARO+Aux would be interpreted as revising the empirical distribution of ERM to a mixture (mixture weight cross-validated) of the empirical and auxiliary distributions. DRO+ARO+Aux, on the other hand, can be interpreted as DRO over a carefully reduced ambiguity set (intersection of the empirical and auxiliary Wasserstein balls). The results are in Table 5. Analogous results follow as before (that is, DRO+ARO+Aux is the ‘winning’ approach, DRO+ARO and ARO+Aux alternate for the ‘second’ approach), with the exception of the dataset contraceptive, where ARO+Aux outperforms others.
@@
12.2 MNIST/EMNIST Experiments
Our setting is analogous to the UCI experiments. However, for auxiliary data, we use the EMNIST dataset. We used the MLDatasets package of Julia to prepare such auxiliary data.
12.3 Artificial Experiments
Data generation

### v4

@@ B.4 Proof of Proposition 2
where the first identity follows from the definition of composition and the second identity employs Toland’s duality (Toland, 1978) to rewrite difference of convex functions optimization.
By using infimal convolutions (Rockafellar, 1997, Theorem 16.4), we can reformulate $g^{*}$:
@@
where the first identity follows from the definition of the convex conjugate, the second identity introduces an additional variable to make this an equality-constrained optimization problem, the third identity takes the Lagrange dual (which is a strong dual since the problem maximizes a concave objective with a single equality constraint), the fourth identity rearranges the expressions, the fifth identity exploits unboundedness of $\bm{x}$, the sixth identity uses the definition of convex conjugates and the final identity replaces the feasible set $\theta\in\mathbb{R}$ with the domain of $L^{\star}$ without loss of generality as this is an $\inf$-problem.
@@
DC maximization terms similar to the one dealt by Lemma 1 appear on the left-hand side of the constraints of Inter-ARO (cf. formulation in Proposition 1). These constraints would admit a tractable reformulation for the case without auxiliary data because the $\inf$-term in the reformulation presented in Lemma 1 does not appear in such cases. To see this, eliminate the second norm (the one associated with auxiliary data) by taking $\widehat{\lambda}=0$, which will cause the constraint $\lvert\theta\rvert\cdot\lVert\bm{z}\rVert_{q^{\star}}\leq\widehat{\lambda}$ to force $\bm{z}=\mathbf{0}$, and the alternative formulation will thus be:
@@
where we used the fact that $L=L^{**}$ and $\sup_{\theta\in\mathrm{dom}(L)}\lvert\theta\rvert=\mathrm{Lip}(L)$ since $L$ is closed convex (Rockafellar, 1997, Corollary 13.3.3). Hence, the DC maximization can be represented with a convex function with an additional convex inequality, making the constraints tractable for the case without auxiliary data. For the case with auxiliary data, however, the $\sup_{\theta}\inf_{\bm{z}}$ structure makes these constraints equivalent to two-stage robust constraints (with uncertain parameter $\theta$ and adjustable variable $\bm{z}$), bringing an adjustable robust optimization (Ben-Tal et al., 2004; Yanıkoğlu et al., 2019) perspective to Inter-ARO. By using the univariate representation $\ell_{\bm{\beta}}^{\alpha}(\bm{x},y)=L^{\alpha}(y\cdot\bm{\beta}^{\top}\bm{x})$Inter-ARO can be written as
@@
This reduces problem (46) to a generic two-stage robust optimization problem with random recourse (Subramanyam et al., 2020, Problem 1) which is proven to be NP-hard even if $S^{\alpha*}$ was constant (Guslitser, 2002). ∎
B.5 Proof of Theorem 1
Consider the reformulation 43 of Inter-ARO that we introduced in the proof of Proposition 2. For any $i\in[N]$ and $j\in[\widehat{N}]$, the corresponding constraint in the first group of ‘adjustable robust’ ($\forall$,\ $\exists$ , ∃) constraints will be:
@@ B.5 Proof of Theorem 1
Notice that this is a safe approximation, since any fixed $\bm{z}$ satisfying the latter system is a feasible static solution in the former system, meaning that for every realization of $\theta$ in the first system, the inner $\exists\bm{z}$ can always ‘play’ the same $\bm{z}$ that is feasible in the latter system (hence the latter is named the static relaxation, Bertsimas et al. 2015). In the relaxed system, we can drop $\forall\theta$ and keep its worst-case realization instead:
@@
The term $\sup_{\theta\in\mathrm{dom}(L^{*})}\{-L^{\alpha*}(\theta)+\theta\cdot\bm{\beta}^{\top}\bm{x}^{i}+\theta\cdot\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i})\}$ is the definition of the biconjugate $L^{\alpha**}(\bm{\beta}^{\top}\bm{x}^{i}+\bm{z}^{\top}(\widehat{\bm{x}}^{j}-\bm{x}^{i}))$. Since $L^{\alpha}$ is a closed convex function, we have $L^{\alpha**}=L^{\alpha}$ (Rockafellar, 1997, Corollary 12.2.1). Moreover, $\sup_{\theta\in\mathrm{dom}(L^{*})}\{\lvert\theta\rvert\}$ is an alternative representation of the Lipschitz constant of the function $L^{\alpha}$ (Rockafellar, 1997, Corollary 13.3.3), which is equal to $1$ as we showed earlier. The adjustable robust constraint thus reduces to:
@@
as a result of the static relaxation. This relaxed reformulation applies to all $i\in[N]$ and $j\in[\widehat{N}]$ as well as to the second group of adjustable robust constraints analogously. Replacing each constraint of 43 with this system concludes the proof.∎
B.6 Proof of Corollary 2
@@
where we also used the definition of $L^{\alpha}$. This problem is identical to DR-ARO, which means that feasible solutions of DR-ARO are feasible for 21 if the additional variables $(\widehat{\lambda},\widehat{\bm{s}},\bm{z}^{l}_{ij})$ are set to zero, concluding the first statement of the corollary.
The second statement is immediate since $\widehat{\varepsilon}\rightarrow\infty$ forces $\widehat{\lambda}=0$ due to the term $\widehat{\varepsilon}\widehat{\lambda}$ in the objective of 21, and this proof shows in such a case 21 reduces to DR-ARO (which is identical to Inter-ARO when $\varepsilon\rightarrow\infty$ by definition). ∎
B.7 Proof of Observation 2
@@ B.8 Proof of Proposition 3
and since $d$ is a feature-label metric (cf. Definition 1) the minimum is well-defined (Villani, 2009, Theorem 4.1). We name the optimal solutions to the above problem the optimal couplings. Let $\Pi^{1}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{1})$ and let $\Pi^{2}$ be an optimal coupling of $\mathrm{W}(\mathbb{Q},\mathbb{P}^{2})$ and define $\Pi^{c}=\lambda\cdot\Pi^{1}+(1-\lambda)\cdot\Pi^{2}$. We have
@@
B.9 Proof of Theorem 2
Since each result in the statement of this theorem is abridged, we will present these results sequentially as separate results. We review the existing literature to characterize $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, in a similar fashion with the results presented in (Selvi et al., 2022, Appendix A) for the logistic loss, by revising them to the adversarial loss whenever necessary. The $N$-fold product distribution of $\mathbb{P}^{0}$ from which the training set $\mathbb{P}_{N}$ is constructed is denoted below by $[\mathbb{P}^{0}]^{N}$.
@@ B.9 Proof of Theorem 2
The statement follows from Theorem 18 of Kuhn et al. (2019). The presented decay rate $\mathcal{O}(N^{-1/n})$ of $\varepsilon$ as $N$ increases is optimal (Fournier and Guillin, 2015). ∎
Now that we gave a confidence for the unknown radius $\varepsilon$ satisfying $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$, we analyze the underlying optimization problems. Most of the theory is well-established for logistic loss function, and in the following we show that similar results follow for the adversarial loss function. For convenience, we state DR-ARO again by using the adversarial loss function as defined in Observation 1:
@@
holds for all $\eta\in(0,1)$ and all optimizers $\bm{\beta}^{\star}$ of 49.
@@
The statement follows from Theorem 19 of Kuhn et al. (2019) given that $\ell^{\alpha}_{\bm{\beta}}$ is a finite-valued continuous loss function. ∎
Theorem 5 states that the optimal value of 49 overestimates the true loss with arbitrarily high confidence $1-\eta$. Despite the desired overestimation of the true loss, we show that 49 is still asymptotically consistent if we restrict the set of admissible $\bm{\beta}$ to a bounded set [^2] Note that, this is without loss of generality given that we can normalize the decision boundary of linear classifiers..
@@
If we show that there exists $\bm{\xi}^{0}\in\Xi$ and $C>0$ such that $\ell^{\alpha}_{\bm{\beta}}(\bm{x},y)\leq C(1+d(\bm{\xi},\bm{\xi}^{0}))$ holds for all $\bm{\beta}\in\mathcal{H}$ and $\bm{\xi}\in\Xi$ (that is, the adversarial loss satisfies a growth condition), the statement will follow immediately from Theorem 20 of (Kuhn et al., 2019).
@@
$\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}+\alpha\cdot\lVert\bm{\beta}\rVert_{p^{\star}}))\leq C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{0}]})$
@@
$C(1+\lVert\bm{x}-\bm{x}^{0}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{0}]})$ $=C(1+\lVert\bm{x}\rVert_{q}+\kappa\cdot\mathds{1}\small{[y\neq y^{0}]})$
@@
So far, we reviewed tight characterizations for $\varepsilon$ so that the ball $\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ includes the true distribution $\mathbb{P}^{0}$ with arbitrarily high confidence, proved that the DRO problem 49 overestimates the true loss, while converging to the true problem asymptotically as the confidence $1-\eta$ increases and the radius $\varepsilon$ decreases simultaneously. Finally, we discuss that for optimal solutions $\bm{\beta}^{\star}$ to 49, there are worst case distributions $\mathbb{Q}^{\star}\in\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ of nature’s problem that are supported on at most $N+1$ outcomes.
@@
If we restrict the hypotheses $\bm{\beta}$ to a bounded set $\mathcal{H}\subseteq\mathbb{R}^{n}$, then there are distributions $\mathbb{Q}^{\star}\in\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ that are supported on at most $N+1$ outcomes and satisfy:
@@
The proof follows from (Yue et al., 2022). ∎
See the proof of Selvi et al. (2022, Theorem 8) and the discussion that follows for insights and further analysis on these results presented.
B.10 Proof of Theorem 3
@@
$\widehat{\varepsilon}_{1}\geq\begin{cases}\left(\dfrac{\log(\widehat{c}_{1}/\eta_{2})}{\widehat{c}_{2}\cdot\widehat{N}}\right)^{1/\max\{n,2\}}&\text{if }\widehat{N}\geq\dfrac{\log(\widehat{c}_{1}/\eta_{2})}{\widehat{c}_{2}}\\ \left(\dfrac{\log(\widehat{c}_{1}/\eta_{2})}{\widehat{c}_{2}\cdot\widehat{N}}\right)^{1/\widehat{a}}&\text{otherwise},\end{cases}$
@@
Appendix C EXPONENTIAL CONIC REFORMULATION OF DR-ARO
@@
Following the conic modeling guidelines of MOSEK ApS (2023), for new variables $v^{+}_{i},w^{+}_{i}\in\mathbb{R}$, the first constraint can be written as
@@
Appendix D FURTHER DETAILS ON NUMERICAL EXPERIMENTS
D.1 UCI Experiments
Preprocessing UCI datasets.
We experiment on 10 UCI datasets (Kelly et al., 2023) (cf. Table 3). We use Python 3 for preprocessing these datasets. Classification problems with more than two classes are converted to binary classification problems (most frequent class/others). For all datasets, numerical features are standardized, the ordinal categorical features are left as they are, and the nominal categorical features are processed via one-hot encoding. As mentioned in the main paper, we obtain auxiliary (synthetic) datasets via SDV, which is also implemented in Python 3.
@@ D.1 UCI Experiments
Detailed misclassification results on the UCI datasets.
Table 4 contains detailed results on the out-of-sample error rates of each method on 10 UCI datasets for classification. All parameters are $5$-fold cross-validated: Wasserstein radii from the grid $\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0,1,2,5,10\}$, $\kappa$ from the grid $\{1,\sqrt{n},n\}$ the weight parameter of ARO+Aux from grid $\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0,1\}$. We fix the norm defining the feature-label metric to the $\ell_{1}$-norm, and test $\ell_{2}$-attacks, but other choices with analogous results are also implemented.
@@
Finally, we demonstrate that our theory, especially DRO+ARO+Aux, contributes to the DRO literature even without adversarial attacks. In this case of $\alpha=0$, ERM and ARO would be equivalent, and DRO+ARO would reduce to the traditional DR LR model (Shafieezadeh-Abadeh et al., 2015). ARO+Aux would be interpreted as revising the empirical distribution of ERM to a mixture (mixture weight cross-validated) of the empirical and auxiliary distributions. DRO+ARO+Aux, on the other hand, can be interpreted as DRO over a carefully reduced ambiguity set (intersection of the empirical and auxiliary Wasserstein balls). The results are in Table 5. Analogous results follow as before (that is, DRO+ARO+Aux is the ‘winning’ approach, DRO+ARO and ARO+Aux alternate for the ‘second’ approach), with the exception of the dataset contraceptive, where ARO+Aux outperforms others.
Cross-validated Wasserstein radii.
In Corollary 2, we discussed the feasibility of ignoring the auxiliary data when its data-generating distribution is distant from the true data-generating distribution. To examine whether large values of $\widehat{\varepsilon}$ are selected via cross-validation in our experiments, we investigated their histograms and observed that this indeed occurs frequently. For example, as shown in Table 5, when there are no adversarial attacks, DRO+ARO and DRO+ARO+Aux yield identical errors on the ‘annealing’ dataset. This is because the two models become equivalent: DRO+ARO+Aux selects a large $\widehat{\varepsilon}$ (either $5$ or $10$), which ensures that the intersection of the Wasserstein balls reduces to the empirical distribution. With $\alpha=0.05$, DRO+ARO+Aux selects a large $\widehat{\varepsilon}$ in 7 out of 10 simulations, while in the remaining 3 it selects a smaller value, resulting in a nontrivial intersection and, in turn, improved performance over DRO+ARO. We conclude that our method selects a nontrivial $\widehat{\varepsilon}$ only when there is evidence that the auxiliary data is useful, that is, when its data-generating distribution is sufficiently close to the true one. For instance, on the ‘absent’ dataset, we always have $\widehat{\varepsilon}\in\{10^{-2},10^{-1}\}$. We also revised our numerical experiments by modifying the Gaussian copula synthesizer to enforce uniform marginals (which results in significant information loss), and in this setting, our models consistently selected large $\widehat{\varepsilon}$ values.
@@
Different training/auxiliary dataset ratio.
Recall that in the UCI experiments, we sampled $15\%$ of the original dataset as the training set, and used $45\%$ to generate a synthetic auxiliary dataset. This setup was chosen to simulate scenarios with limited training data, where overfitting is a particular concern. If we reverse these proportions and use $45\%$ of the original dataset for training and $15\%$ for generating auxiliary data instead, we find that the best-performing method remains DRO+ARO+Aux, and the relative ranking among the models remains qualitatively similar. One notable difference is that the improvement of the auxiliary-data-oblivious DRO+ARO method over the non-DR ARO+Aux model becomes less significant, and is even reversed in the case of $\alpha=0.05$ on the ‘ecoli’ dataset: ERM ($14.59\%$), ARO ($11.41\%$), ARO+Aux ($9.04\%$), DRO+ARO ($9.41\%$), DRO+ARO+Aux ($8.89\%$). As expected, since more data is drawn from the true data-generating distribution, all methods exhibit improved out-of-sample performance compared to the original setup. However, in this case, DRO+ARO no longer outperforms ARO+Aux.
D.2 MNIST/EMNIST Experiments
The setting in the MNIST/EMNIST experiments is similar to that in the UCI experiments. However, for auxiliary data, we use the EMNIST dataset which we accessed via the MLDatasets package of Julia.
Moreover, in §2 we reviewed the literature showing that when statistical error is not a concern, that is, when optimizing over the empirical distribution cannot cause overfitting (e.g., in high-data regimes), then adversarial training is equivalent to a type-$\infty$ Wasserstein DRO problem with radius $\varepsilon=\alpha$. Hence, a natural question is whether increasing the value of $\varepsilon$ further also provides distributional robustness. To this end, in Table 6, we revise Table 2 and add an additional benchmark DRO++. Here, we take $\varepsilon=\alpha+\varepsilon^{\prime}$, and cross-validate $\varepsilon^{\prime}$ from the same grid that we cross-validated $\varepsilon$ for methods DRO+ARO and DRO+ARO+Aux. We observe that DRO++ does not improve over DRO+ARO, which is expected given that type-$\infty$ Wasserstein DRO does not provide better generalizations, unlike type-$1$ Wasserstein DRO (cf. the discussion at the end of §3). Yet, this method improves over ARO in all cases, and even over ARO+Aux in the no-attack or $\ell_{\infty}$-attack settings.
Table 6: Additional MNIST/EMNIST Benchmark.
Attack ERM ARO ARO+Aux DRO++ DRO+ARO DRO+ARO+Aux
No attack ($\alpha=0$) 101.55% 101.55% 101.19% 100.72% 100.64% 100.53%
$\ell_{1}$ ($\alpha=68/255$) 102.17% 101.84% 101.33% 101.40% 100.66% 100.57%
$\ell_{2}$ ($\alpha=128/255$) 199.93% 103.36% 102.54% 102.72% 102.40% 102.12%
$\ell_{\infty}$ ($\alpha=8/255$) 100.00% 102.60% 102.38% 102.31% 102.20% 101.95%
D.3 Artificial Experiments
Data generation.